How many moles of NaOH are there in 50cm^3 of a solution that has a molarity of 1.76?
1
Expert's answer
2019-07-31T08:16:03-0400

Given:

"V(NaOH) = 50cm^3 = 50mL = 0.05L"

"c(NaOH)=1.76 \frac{mol}{L}"

According to the equation of molar concentration, "c(NaOH)=\frac{n(NaOH)}{V(NaOH)}"

Thus, amount of NaOH in moles:

"n(NaOH)=c(NaOH) \cdot V(NaOH) = 1.76\frac{mol}{L}\cdot0.05L=0.088mol"